OpenStudy (anonymous):

a man is in a racing area instead of counting the number of humans and horses there, he instead counted 74 heads and 196 legs, however the man knew the number of humans and hoarses, how did he do it? how many people and hoarse are there????

OpenStudy (anonymous):

?

OpenStudy (anonymous):

before we continue what are your credentials?? if you dont mind me asking

OpenStudy (dumbcow):

Let A be number of humans, B be number of horses human has 2 legs, horse has 4 legs A+B = 74 2A+4B = 196 solve system using substitution or elimination ... credentials? college degree in math

OpenStudy (anonymous):

so to be clear how man humans and how many horses?

OpenStudy (dumbcow):

you have to solve system of 2 equations with 2 unknowns Multiply top equation by -2, then add the equation to eliminate variable A -2A -2B = -148 2A + 4B = 196 ----------------- 2B = 48 Solve this for B, then substitute back in to find A can you take it from here?

OpenStudy (anonymous):

id like you to continue ma'am/or sir!

OpenStudy (dumbcow):

B = 48/2 B = 24 A+B = 74 A+24 = 74 A = 74-24 = 50 There are 50 humans and 24 horses

OpenStudy (anonymous):

you happen to be right b/c i was also computing the question as well

OpenStudy (anonymous):

in your honest opinion was this question hard for you?

OpenStudy (dumbcow):

sounds like you are taking a survey or something? do you really need help in math?? no this was not hard for me, but im here to help not ask questions